# Fostering Collective Discourse: A Distributed Role-Based Approach to Online News Commenting

### v1

@@ 5.2.4. Politeness
Lastly, to examine effects on politeness, we extracted sentence-level strategies using ConvoKit and grouped them into positive politeness (e.g., please, gratitude, apology) and negative or impolite strategies (e.g., direct address, swearing, negation). For each comment, a politeness score was computed, and participant-level averages were analyzed across conditions using mixed-effects linear models with participant random intercepts and article variance components. Results indicated a baseline mean of 0.573 and a system mean of 0.581 ($\Delta$ = +0.007; $\beta$ = 0.007, SE = 0.009, 95% CI [-0.011, 0.025]), showing no significant difference. These findings suggest that politeness levels remained stable, with the system neither enhancing nor diminishing prosocial tone, thereby preserving civility across conditions.
@@
The post-survey results showed that participants responded positively to the system’s ability to improve their understanding of issues, discussion flow, and diverse viewpoints. Guided topics and questions were particularly effective in helping participants identify key aspects of the articles, stay focused on central points without being distracted by large volumes of comments, and articulate their own thoughts more clearly. Clustering and summarization features provided concise overviews of discussion threads, allowing participants to grasp core ideas quickly, spot gaps, and connect with related opinions. These features also broadened perspectives by grouping similar but differently worded comments and highlighting important flows of discussion, which encouraged participants to think more fluently and from multiple angles. Overall, participants indicated that the system enhanced their comprehension and engagement across three stages of commenting: reading and understanding, structuring and writing, and engaging meaningful discussion. The following sections present the analyzed results of interviews about how our system supported each stage of this process.
@@ 6. Discussion
Our system demonstrated that fostering collective aspects of user participation can lead to more constructive and meaningful discussions, significantly enhancing the overall quality of discourse. Throughout the study, we found that this improvement was driven by two key factors: 1) the implementation of distributed roles, and 2) the impact of the collective output within a structured discussion space.
@@ 6.1.1. Hierarchical Structure and Aggregated Viewpoints
Our system’s hierarchical view of discussion – incorporating threads, summarization, clusters, and comments – was designed to offer a structured output, enabling users to navigate and streamline their focus efficiently. However, it also risked overlooking important details and nuanced individual perspectives. For instance, P11 highlighted the significance of precise personal viewpoints in comments, particularly in discussions about serious and important issues, suggesting that summarization might sometimes obscure critical insights if not carefully managed. This underscores the design decisions of generating views that balance aggregated opinions with the retention of essential details.
@@ 6.2. Navigating Trade-off Values in Collective Discourse
Building collective discourse allowed users to explore diverse viewpoints, fostering a deeper understanding of complex issues. However, maintaining such discussions imposes significant workload on users tasked with organizing and structuring the conversation, which can divert their focus from engagement in commenting. In real-world settings, further efforts should focus on leveraging the collective effort and scale of the community to distribute the workload more effectively. This could involve designing smaller, more manageable tasks, such as tagging or summarizing portions of the discussion. Additionally, systems could actively incentivize users who are willing to take on more responsibility, recognizing that not all users contribute equally. By focusing the workload on those willing and able to contribute more, we can create a more sustainable approach to managing collective discourse.
@@
Our findings showed that the system brought like-minded individuals together, which positively impacted users by helping them connect with others who shared similar perspectives and reduced the burden of expressing their opinions. However, this also raises concerns about the potential risk of echo chambers, as echo chambers can emerge when groups form around shared views, reinforcing their beliefs while excluding opposing opinions (Cinelli et al., 2021). Interestingly, our findings indicate that bringing like-minded individuals together did not necessarily lead to reduced diversity or increased polarization, as different steps in the commenting process actively supported both aspects. While the system provided common ground for easier engagement during the contribution phase, it also simultaneously offered overviews of diverse viewpoints and reminders of overlooked perspectives during the reading and structuring phases. By thoughtfully balancing these aspects, we highlight the potential for creating discussion environments that maintain diversity without sacrificing the benefits of shared focus.
@@
6.3. Future Work
@@ 6.4. Limitations
While our study provides valuable insights into designing collaborative commenting systems, several limitations should be acknowledged. One limitation is the relatively small number of participants involved in the study, with 38 users interacting with our system. In larger discussion spaces, uncivil behaviors can become more prevalent and problematic, impacting a broader audience. A larger participant pool could provide a deeper understanding of how the user performs their actions in larger-scale discussions and the system’s ability to manage a higher volume of comments while maintaining coherence in extended threads.
Beyond the scale of participants, the duration of the study also posed a limitation. Our study was conducted over a relatively short period–three days per condition, with one day allocated for each article–whereas participation in real-world commenting systems is not typically time-restricted. Although our one-day participation design for each article was informed by the observation that articles often receive the highest volume of attention upon initial publication (Kaltenbrunner et al., 2007), future research should consider extending the study duration to enable a more comprehensive, long-term analysis of commenting behavior.
In addition, to create a controlled experimental environment, we pre-assigned user levels to each participant, even though this approach is not directly applicable in real-world settings. Alternative designs could allow user levels to be adjusted dynamically, such as through a level-up system based on the completion of specific number of activities or by aligning with user preferences. However, in a short-term study where participants engaged with each article for a single day, maintaining a fixed number of participants per level made it impractical for users to experience role changes based on their level of contribution. For future long-term studies, alternative strategies for role assignment could be explored to better reflect real-world dynamics and power structure of role hierarchies.
Another limitation is the restricted range of news sources used in the study, which limited the evaluation to a narrow set of articles and topics, potentially missing the full spectrum of content and perspectives in broader discussions. Future research should include a wider variety of topics and additional news sources to provide a more comprehensive assessment of the system’s effectiveness. Additionally, the majority of participants were young college students from university communities, which may limit the applicability of the findings to other demographics. Including a more diverse participant pool in future studies will help capture a broader range of user interaction patterns and motivations.
@@ A.0.2. Suggested Thread Topics and Guiding Questions
Article text: <text>

### v2

@@ 5.2.4. Politeness
Lastly, to examine effects on politeness, we extracted sentence-level strategies using ConvoKit and grouped them into positive politeness (e.g., please, gratitude, apology) and negative or impolite strategies (e.g., direct address, swearing, negation). For each comment, a politeness score was computed, and participant-level averages were analyzed across conditions using mixed-effects linear models with participant random intercepts and article variance components. Results indicated a baseline mean of 0.573 and a system mean of 0.581 ($\Delta$ = +0.007; $\beta$ = 0.007, SE = 0.009, 95% CI [-0.011, 0.025]), showing no significant difference. These findings suggest that politeness levels remained stable, with the system neither enhancing nor diminishing prosocial tone, thereby preserving civility across conditions. Accordingly, these results do not support H5, indicating that structured participation preserved politeness in discussion.
@@
The post-survey results showed that participants responded positively to the system’s ability to improve their understanding of issues, discussion flow, and diverse viewpoints. Guided topics and questions were particularly effective in helping participants identify key aspects of the articles, stay focused on central points without being distracted by large volumes of comments, and articulate their own thoughts more clearly. Clustering and summarization features provided concise overviews of discussion threads, allowing participants to grasp core ideas quickly, spot gaps, and connect with related opinions. These features also broadened perspectives by grouping similar but differently worded comments and highlighting important flows of discussion, which encouraged participants to think more fluently and from multiple angles. Overall, participants indicated that the system enhanced their comprehension and engagement across three stages of commenting: reading and understanding, structuring and writing, and engaging in meaningful discussion. The following sections present the analyzed results of interviews about how our system supported each stage of this process.
@@ 6. Discussion
Our system demonstrated that fostering collective aspects of user participation can lead to more constructive and meaningful discussions, significantly enhancing the overall quality of discourse. Throughout the study, we found that this improvement was driven by two key factors: (1) the implementation of distributed roles, and (2) the impact of the collective output within a structured discussion space.
@@ 6.1.1. Hierarchical Structure and Aggregated Viewpoints
Our system’s hierarchical view of discussion – incorporating threads, summarization, clusters, and comments – was designed to offer a structured output, enabling users to navigate and streamline their focus efficiently. However, it also risked overlooking important details and nuanced individual perspectives. For instance, P11 highlighted the significance of precise personal viewpoints in comments, particularly in discussions about serious and important issues, suggesting that summarization might sometimes obscure critical insights if not carefully managed. Although participants did not report distortions or unnecessary summarization during the study, the findings underscore the importance of ensuring that collaborative structures make the underlying reasoning not only streamlined but also transparent and recoverable for users who seek deeper engagement.
@@ 6.2. Navigating Trade-off Values in Collective Discourse
Building collective discourse allowed users to explore diverse viewpoints, fostering a deeper understanding of complex issues. However, maintaining such discussions imposes a significant workload on users tasked with organizing and structuring the conversation, which can divert their focus from engagement in commenting. In real-world settings, further efforts should focus on leveraging the collective effort and scale of the community to distribute the workload more effectively. This could involve designing smaller, more manageable tasks, such as tagging or summarizing portions of the discussion. Additionally, systems could actively incentivize users who are willing to take on more responsibility, recognizing that not all users contribute equally. By focusing the workload on those willing and able to contribute more, we can create a more sustainable approach to managing collective discourse.
@@
Our findings showed that the system brought like-minded individuals together, which positively impacted users by helping them connect with others who shared similar perspectives and reducing the burden of expressing their opinions. However, this also raises concerns about the potential risk of echo chambers, as echo chambers can emerge when groups form around shared views, reinforcing their beliefs while excluding opposing opinions (Cinelli et al., 2021). Interestingly, our findings indicate that bringing like-minded individuals together did not necessarily lead to reduced diversity or increased polarization, as different steps in the commenting process actively supported both aspects. While the system provided common ground for easier engagement during the contribution phase, it also simultaneously offered overviews of diverse viewpoints and reminders of overlooked perspectives during the reading and structuring phases. By thoughtfully balancing these aspects, we highlight the potential for creating discussion environments that maintain diversity without sacrificing the benefits of shared focus.
@@
6.3. Implications for Social Dynamics in Online Discussions
While our system demonstrated the potential of structured collaboration in online news discussion, reorganizing comments fundamentally reshapes the social dynamics through which discourse emerges. We discuss implications in relation to user motivation, social influence patterns, and sustainable participation at scale.
6.3.1. Supporting Diverse Motivations for Participation
@@
6.3.2. Managing Conformity and Social Influence
Re-structuring comments also restructures how social cues are revealed, which may shape influence and conformity behaviors. Prior works in social computing demonstrate that individuals often rely on visible popularity cues and majority viewpoints when forming opinions (Wijenayake et al., 2020; Colliander, 2019). By design, our system counteracts some of these effects by presenting a more evenly weighted distribution of perspectives and reducing early dominance patterns typically reinforced by chronological or popularity-driven feeds. However, aggregated summaries, if not carefully contextualized, risk being perceived as consensus rather than as one representation of the discussion. As a result, while fragmentation and extreme polarization may be reduced, there remains a possibility that conformity pressures become subtler, guiding users toward what appears collectively endorsed and diminishing opportunities for less prominent viewpoints to surface. To ensure that collaboration supports deliberation rather than conformity, future system designs should consider mechanisms that make the diversity and provenance of viewpoints legible, allowing users to appreciate both common ground and unresolved disagreement.
6.3.3. Incentivizing Knowledge-Building at Scale
A collaborative system introduces additional coordination effort, which becomes particularly challenging in real-world situations where discussions unfold rapidly or reach large scale. As open collaboration scales, increased governance needs can burden or discourage newcomers (Halfaker et al., 2013), and the organizational labor that sustains communities often remains invisible and undervalued (Matias, 2019b), challenging the long-term sustainability of user-driven knowledge building. Participants in our study welcomed the sense of agency associated with curating and organizing discourse, yet in settings with hundreds or thousands of comments especially during breaking news events, the cognitive burden of coordination could outweigh the perceived benefits. The sustainability of such systems thus depends not only on interaction design but also on how they support meaningful labor. This may involve intelligent automation that reduces repetitive tasks, lightweight contribution opportunities that allow users to help without extensive time investment, and incentive structures that recognize and reinforce knowledge-building behaviors over passive or reactive participation. Rather than minimizing effort entirely, the goal is to align users’ invested effort with visible improvements in collective understanding and discourse quality.
@@ 6.4. Limitations
While our study provides valuable insights into designing collaborative commenting systems, several limitations should be acknowledged to contextualize our findings and guide future work.
6.4.1. Methodological Scale and Sampling Constraints
First, the study involved a relatively small number of participants (N=38) interacting within a controlled discussion setting. In larger discussion spaces, uncivil behaviors can become more prevalent and problematic, impacting a broader audience. Future work should evaluate the system’s performance at scale to understand its robustness in maintaining coherence within extended and high-velocity threads.
Additionally, the study was conducted over a comparatively short duration – three days per condition with one article per day. Although this design reflects common attention patterns in real-world news consumption (Kaltenbrunner et al., 2007), longer-term deployments are needed to observe how participation evolves over sustained engagement cycles and whether collaboration dynamics persist beyond initial novelty.
The demographic composition of our participants also limits generalizability. Most users were university students with high digital literacy, whereas real-world platforms often host a broader range of ages, backgrounds, and commenting norms. Moreover, as all interactions occurred in English, varying levels of language proficiency may have shaped how comfortably participants articulated arguments or engaged in coordination. Future studies should include more diverse populations and content sources to capture a wider variety of interaction styles, language abilities, and topic domains.
6.4.2. Limited Ecological Representation of Role Hierarchies
To create a controlled environment for examining role-based collaboration, users were assigned fixed levels rather than earning roles dynamically. While this allowed systematic observation of how roles influence participation, it does not fully represent how power structures form and shift in real-world communities. Alternative designs could allow user levels to be adjusted dynamically, such as through a level-up system based on the completion of specific number of activities or by aligning with user preferences. Longer engagements could incorporate adaptive role mechanisms to better reflect evolving power structures and authority within online discourse.
6.4.3. Design Limitations in Supporting Constructive Depth and Novelty
Despite improvements in perspective balance and user participation, the system showed limitations in supporting explicitly reasoned and novelty of contributions. The observed reduction in Supported-Claim Ratio (SCR) suggests that, while users remained engaged, their comments contained fewer explicit justifications. One plausible interpretation is that the structured elements of the system – such as guided prompts and cluster-level summaries – captured much of the shared reasoning, reducing the need for users to restate evidence at the level of individual comments. Although this may reduce redundancy, it can obscure the transparency and traceability of argumentative depth. Ensuring that collective structures remain accountable to the arguments they represent will therefore require interface mechanisms that make the underlying rationale easily discoverable when needed.
Additionally, although the system successfully prevented the dominance of majority viewpoints by balancing visibility (H_norm), it did not significantly expand the overall range of distinct viewpoints (Coverage). This suggests that collaboration may help users converge on and refine existing ideas rather than encouraging the emergence of entirely new ones. While convergence can be beneficial for coherence and shared understanding, constructive discourse also relies on the introduction of novel or previously underrepresented insights. Designing features that support viewpoint evolution – such as prompts for alternative reasoning, mechanisms for merging across clusters, or identifying missing angles – may help prevent discussions from stabilizing prematurely around familiar views.
@@ 7. Conclusion
Acknowledgements.
This research was supported by the KAIST New Faculty Research Fund (Project No. G04230040: Building Integrated Models for Online Governance).
@@ A.0.2. Suggested Thread Topics and Guiding Questions
Article text: { text }